You invest $12,000 in two funds paying 5.5% and 7.5% simple interest. The total annual interest is $772.50. How much is invested in each fund?

Answer by checkley71(8403)

.075*X+.055(12,000-X)=772.50
.075X+660-.055X=772.50
.02X=772.50-660
.02X=112.50
X=112.50/.02
X=5,625 IS THE AMOUNT INVESTED @ 7.5%.
THUS12,000-5,625=6,375 INVESTED @ 5.5%
PROOF:
.075*5625+.055*6,375=772.50
421.875+350.625=772.50
772.50=772.50